OpenStudy (anonymous):

There are five more children in Miss Brown's class than in Mr whites class. Ten children move from MrS.Browns class to mr. Whites class. now there are twice as many children in Mr. Whites as in Mrs. Browns room. How many children were in Miss browns room before the move?

OpenStudy (anonymous):

We set up a system of equations B=W+5 (5 more in Brown than White) 2(B-10)=W+10 (move 10 from B to W and W has twice what B has) Now plug B-5=W into the second equation and solve for B. 2(B-10)=B+5 B=25. W=20 so we can check B-10=15, W+10=30. 15*2=30. So we are correct.

OpenStudy (anonymous):

does it make sense?